[Hallc_running] [New Logentry] RC report

arunts at jlab.org arunts at jlab.org
Fri Nov 12 09:05:02 EST 2021

Logentry Text:
After some intermittent beam issues during yesterday's day shift we had smooth swing and owl shifts with good quality beam. After collecting enough H(e,e')ps events (page 4 of the run plan), the spectrometers were rotated (https://logbooks.jlab.org/entry/3940680, https://logbooks.jlab.org/entry/3940689) to start taking H(e,e')p check coincidence runs (page 5 & 6 of the run plan). At around on 3:20 PM radcon, Steve Lassiter & co, Jerry & co went downstairs for a monitored small angle and minimum HMS+SHMS opening angle rotation (https://logbooks.jlab.org/entry/3940924). The rotation went quite smoothly with HMS moved to 11.01 degrees and SHMS moved to within 18.00 degrees of the HMS or 6.99 degrees (https://logbooks.jlab.org/entry/3941032). After HMS Q2 Interlock issues were fixed (thanks, Ellen!) we continued data taking through the rest of the swing and owl shifts (pages 8 - 13 of the run plan).

Some highlights from yesterday and early morning’s running:
1. SHMS and HMS small angle rotation (https://logbooks.jlab.org/entry/3940105).
2. Ellen's access to fix HMS Q2 interlock issues (https://logbooks.jlab.org/entry/3941171).
3. Downward drift of the beam energy (https://logbooks.jlab.org/entry/3941048).
4. Dave Gaskell's study on good Heep coin events (https://logbooks.jlab.org/entry/3940885).
5. go_cameras updated by Brad Sawatzky (https://logbooks.jlab.org/entry/3941075).

Let's have an RC meeting at 10:00 AM EST
Connection information: https://bluejeans.com/861439920


This is a plain text email for clients that cannot display HTML.  The full logentry can be found online at https://logbooks.jlab.org/entry/3941477
-------------- next part --------------
An HTML attachment was scrubbed...
URL: <https://mailman.jlab.org/pipermail/hallc_running/attachments/20211112/66e4cb84/attachment.html>

More information about the Hallc_running mailing list